OpenStudy (anonymous):

Solve the following system of equations: x + 4y - z = 6 2x - y + z = 3 3x + 2y + 3z = 16

OpenStudy (squirrels):

Do you prefer the substitution method or elimination? or does it not matter.

OpenStudy (solomonzelman):

matrix ?

OpenStudy (anonymous):

Elimination, but it doesn't really matter.

OpenStudy (solomonzelman):

it really looks like a matrix problem.

OpenStudy (anonymous):

What do you mean matrix problem? I need to solve for x y z.

OpenStudy (anonymous):

Yep. I'm in precalc. 10th Grade.

OpenStudy (solomonzelman):

Well, I am saying that I would prefer matrix, but you don't really have to use it if you don't like it.

OpenStudy (anonymous):

How would you do that?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \left[\begin{matrix}? & ? & ?~~~|~~~? \\ ? & ? & ?~~~|~~~? \\ ? & ? & ?~~~|~~~? \end{matrix}\right] }\) on the right on each row, after the " | " you have the constant that the equation is equivalent to, and inside you will in the variable coefficients. \(\large\color{black}{ \displaystyle x + 4y - z = 6 \\ 2x - y + z = 3 \\ 3x + 2y + 3z = 16 }\) \(\large\color{black}{ \displaystyle \left[\begin{matrix}1 & 4 & -1~~~|~~~6 \\ 2 & -1 & 1~~~|~~~3 \\ 3 & 2 & 3~~~|~~~16 \end{matrix}\right] }\)

OpenStudy (solomonzelman):

BUT, this is only I am allowed to do when variables in each row correspond to appropriate column. like first column (if you look at 3 equations) is X's, 2nd column is Y's and etc...

OpenStudy (solomonzelman):

then by manipulating row (multiplying adding subtracting and all that) we must get to \(\large\color{black}{ \displaystyle \left[\begin{matrix}1 & 0 & 0~~~|~~~a \\ 0 & 1 & 0~~~|~~~b \\ 0 & 0 & 1~~~|~~~c \end{matrix}\right] }\) where as we are done, we will have x=a x=b x=c (for whatever numbers you have as your a b c )

OpenStudy (solomonzelman):

if you haven't done those before though, lets use some more usual approaches.

OpenStudy (anonymous):

How do you manipulate the rows to get there?

OpenStudy (solomonzelman):

I mean x=a, y=b, z=3 excuse me

OpenStudy (solomonzelman):

by adding and subtracting row... just the same way as by doing so with regular equations.

OpenStudy (solomonzelman):

except that this is at times more convenient when you have a lot of variables. but again, if this is not something you are well familiar with, then lets go with other approaches (like substitution).

OpenStudy (anonymous):

How do you do that with 3?

OpenStudy (solomonzelman):

how to do what with 3 variables ?

OpenStudy (solomonzelman):

the substitution you will need to sub into the other 2 (remaining) equations.

OpenStudy (anonymous):

3 equations

OpenStudy (solomonzelman):

x + 4y - z = 6 2x - y + z = 3 3x + 2y + 3z = 16 rearranging the first equation for z x + 4y - z = 6 x + 4y = 6 + z `x + 4y - 6 = z` so the new system we will use is: x + 4y - 6 = z 2x - y + z = 3 3x + 2y + 3z = 16 NOW, the substitution comes in handy. 2x - y + (x + 4y - 6) = 3 3x + 2y + 3(x + 4y - 6) = 16

OpenStudy (solomonzelman):

then you have a regular system of equations. After you find the x and y, substitute the values of x and y (that you have found) into any of the ORIGINAL (from very beginning) equations to find z.

OpenStudy (anonymous):

So you just sub in one equation into the others.